OpenStudy (anonymous):

calculate the area of the circle 10 in round your answer to the nearest hundredth

OpenStudy (3psilon):

10in as in 10 inches?

OpenStudy (anonymous):

yes

OpenStudy (3psilon):

so the circle is 10in round which is the circumference

OpenStudy (3psilon):

the equation for circumference is 2*pi*r so 10 = 2*pi*r

OpenStudy (3psilon):

So we need to solve for r divide both sides by 2pi and we get 10/2pi=r r = 1.6~ using a calculator

OpenStudy (3psilon):

Now we plug r into the area of a circle formula which is pi*r^2

OpenStudy (3psilon):

3.14*(1.6)^2= 8

OpenStudy (3psilon):

so the area of the circle is 8in

OpenStudy (3psilon):

about 8.04 when rounding to the nearest hundredth